# refs.bib {refsection}

### v2

@@ 3.5 MedVAL Performance Ablation
Our ablation study (Supplementary Figure 6) investigates the source of MedVAL’s performance improvements, demonstrating that gains arise from the "combined effect" of leveraging distillation via a frontier LM (GPT-4o) and MedVAL’s filtering strategy. We examine this in two parts: 1) whether performance improves even when the LM self-generates synthetic data, and 2) the role of MedVAL’s filtering mechanism in overall performance.
First, we explore whether gains persist when LMs are fine-tuned using self-distillation, without leveraging GPT-4o for synthetic data generation. If performance improves under this setting, it would suggest that MedVAL’s strategy contributes independently of the teacher model’s strength. We observe performance gains even when the LM itself is leveraged as the teacher alongside data filtering ($\mathcal{M}_{\text{MedVAL}}\geq 0.9$). For instance, Llama-3.2-3B’s F1 score increased from 12.8% to 22.1% when using itself as the teacher. Gains were also seen in Qwen3-4B (42.8% $\to$ 43.3%), Llama-3.1-8B (25.9% $\to$ 29.1%), GPT-4o Mini (47.4% $\to$ 50.4%), and GPT-4o (54.5% $\to$ 58.7%), indicating that the MedVAL framework contributes to performance improvements.
Next, we isolate the contribution of filtering by comparing fine-tuning on filtered versus unfiltered synthetic data. When applying the original filtering threshold $\mathcal{M}_{\text{MedVAL}}\geq 0.9$, 57% of the data (1,131 examples) is retained from the original synthetic dataset of 2,000 samples. In contrast, using a relaxed threshold $\mathcal{M}_{\text{MedVAL}}\geq 0.0$ retains all 2,000 training samples, allowing us to also test whether additional unfiltered data necessarily leads to better performance. We find that without filtering, performance only marginally improves or even declines, compared to the relatively larger improvements seen when applying the original filtering threshold. For example, Qwen3-4B sees a drop in F1 score from 42.8% to 39.7%, and Llama-3.2-3B and Llama-3.1-8B see minor improvements from 12.8% to 13.6% and 25.9% to 28.3%, respectively. These results suggest that eliminating low-quality training examples contributes to MedVAL’s performance gains.
@@ 4 Discussion
We introduce MedVAL, a generalizable, self-supervised framework for training LMs to validate LM-generated medical text following a physician-defined taxonomy of risk levels and error categories. By combining synthetic data generation, data filtering, and fine-tuning, MedVAL improves a LM’s baseline capacity across diverse medical tasks. We validate MedVAL on MedVAL-Bench, a benchmark of 840 physician-annotated examples spanning 6 medical tasks. MedVAL extends prior methods by providing a framework that: 1) trains scalable evaluators without physician-in-loop supervision, 2) assesses medical text in the absence of reference outputs or retrieval, 3) supports multilingual evaluation, and 4) offers interpretable, expert-aligned assessments.
@@
Across all settings, MedVAL improved average four-class baseline F1 scores from 36.7% to 51.0%, with MedVAL GPT-4o achieving the highest score on the benchmark. MedVAL also improved average safe/unsafe binary F1 scores from 66.2% to 82.8%, indicating reliable discrimination between safe and unsafe responses. Risk-level analysis revealed that MedVAL enhances model sensitivity, particularly at intermediate risk levels (2–3), which are critical for deciding human review. Task-wise results confirmed strong generalization across in-distribution and out-of-distribution settings, with up to 84% relative improvement on held-out tasks. Notably, MedVAL displayed strong improvements on dialogue2note, the longest input context (average $\sim 1.5$k tokens) out-of-distribution task, demonstrating robustness on challenging, real-world medical tasks.
@@
Similar to MedVAL, there exist automated methods for evaluating LM-generated medical text that have surfaced. However, these methods are either not medical domain-specific or require physicians-in-the-loop, the availability of reference outputs, or external knowledge bases for retrieval. MedVAL does not require either of these constraints. FActScore min2023factscore uses domain-specific extractors for checking claims, though it relies on structured knowledge bases. AlignScore zha2023alignscore proposes a unified alignment function for factuality assessment across tasks such as NLI, QA, and summarization. These methods, including model confidence calibration approaches wang2024calibrating; band2024linguistic; liu2024enhancing, often lack the necessary nuance for clinically focused error assessment. In medical settings, there is often a continuum of right and wrong that requires a very nuanced understanding, making it uniquely challenging compared to general domains. Several studies have focused on error detection for LM-generated medical text. MedHAL mehenni2025medhal, a benchmark for hallucination detection in medical text, relies on physician error annotations limited to specific curated medical tasks. DocLens xie2023doclens introduces multi-aspect metrics—completeness, conciseness, and attribution—tailored for medical text generation. While they report higher agreement with physician assessments than existing metrics, their evaluation relies on the availability of reference outputs. VeriFact chung2025verifact uses retrieval-based evidence to verify statements and is oriented towards multi-document summarization without the capability to train LMs. Recent studies have also focused on error assessment for text generation in radiology: ReXTrust hardy2024rextrust presents a fine-grained hallucination detector, ReXErr rao2024rexerr injects clinical errors into radiology reports, GREEN ostmeier2024green identifies significant errors in radiology reports, and FineRadScore huang2024fineradscore evaluates radiology report generation. While these efforts are crucial for identifying errors, their focus on radiology tasks (e.g., chest X-rays zambrano2025clinically; yu2023evaluating) and reliance on reference outputs limits their generalizability. A detailed comparison of these methods is summarized in Supplementary Table 4.
@@
Despite MedVAL’s broad applicability, our study has limitations. First, the prompt design for generating outputs and filtering thresholds was manual, and more systematic prompt engineering or hyperparameter tuning may further optimize performance. Second, while our input-based validation approach enables evaluation without references, this can limit effectiveness in tasks where the input lacks sufficient information for verification, such as question-answering (medication2answer). Nevertheless, if reference outputs are available, they can be provided as an input to MedVAL for a more holistic assessment, highlighting a beneficial extension of our framework. Furthermore, for QA tasks, we aimed to explore whether safety-critical decisions can be reliably performed, regardless of whether the knowledge originates from the input or the LM’s knowledge base. Our framework demonstrated improved alignment with physicians in the medication2answer task, suggesting that MedVAL can enhance an LM’s ability to assess outputs leveraging its knowledge base. Third, our study framework in terms of model selection and medical tasks was not exhaustive. The six medical tasks we chose may not encompass the full spectrum of clinical document types and scenarios. Finally, while we selected representative open-source and proprietary models for fine-tuning under practical resource settings, results may differ with larger models or other closed platforms. However, we open-source our code and trained models to allow the community to continually validate on broader tasks and models of interest.
@@ 5 Conclusion
In this research, we introduce MedVAL, a novel self-supervised framework for training LMs to validate LM-generated medical text. Our comprehensive evaluation highlights substantial improvements in model performance, achieving better alignment with physicians. The MedVAL-Bench dataset, with its diverse tasks and physician annotations, establishes a robust benchmark for validating LM-generated medical text. To support further research and adoption, we openly release: 1) our codebase, 2) the MedVAL-Bench dataset, and 3) MedVAL-4B, the best-performing open-source model. Together, these contributions lay a foundation for developing and deploying safer, more trustworthy tools to automate the process of validating LM-generated medical text, reducing the burden on physicians. The findings from this study suggest that incorporating MedVAL into clinical workflows could enhance the reliability of medical text generation, thereby improving patient safety and clinical efficiency. Future research should explore prospective studies to validate the practical benefits of MedVAL in real-world settings, potentially leading to an integrated approach that can support healthcare professionals and enhance patient care.
@@ 6 Data and Code Availability
Datasets used in this study include MedicationQA medicationqa, MeQSum MeQSum, Open-i demner2016preparing, MIMIC-IV johnson2020mimic, MIMIC-IV-BHC aali2024mimic and ACI-Bench yim2023aci; abacha2023overview; MEDIQA-Sum2023. All these datasets (except MIMIC) are fully open-source. The MIMIC-IV datasets require PhysioNet access and are currently undergoing review to be made available on PhysioNet. Further distribution of datasets is subject to the data sharing agreements stipulated by the original creators.
@@ 8 Author Contributions
[…]
Table 4: MedVAL comparison with related methods that focus on evaluation of LM-generated text. Here, BHC refers to "brief hospital course".
Method Focus Medical domain Train -able Physician -free training Reference -free evaluation Retrieval -free evaluation Multi -lingual evaluation
FActScore General
AlignScore General
FineRadScore Radiology
ReXTrust Radiology
GREEN Radiology
VeriFact BHC
DocLens Medical
MedHAL Medical
MedVAL Medical
Algorithm 1 MedVAL self-supervised training
1:Frozen generator $g_{\theta}$, frozen validator $v_{\phi}$, fine-tunable validator $v_{\alpha}$, inputs $\mathcal{D}=\{x_{i}\}$, threshold $\tau$
2:Trained validator $v^{*}_{\alpha}$
3:Initialize training dataset $\mathcal{D}_{\text{train}}\leftarrow\emptyset$
4:for $x\in\mathcal{D}$ do
5:     $\delta\leftarrow\textit{RandomChoice}(\{\delta_{1},\delta_{2},\ldots,\delta_{L}\}\mid\delta\in[0,1])$
6:     $\hat{y}\leftarrow y\text{ if available, else }g_{\theta}(x)$triangleright Unperturbed output
7:     $\hat{y}_{\delta}\leftarrow g_{\theta}(x_{\delta})$triangleright Perturbed output
8:     $\hat{\delta}_{\text{clean}}\leftarrow v_{\phi}(x,\hat{y})$triangleright Factual degradation of $\hat{y}$ in comparison to $x$
9:     $\hat{\delta}_{\text{corrupt}}\leftarrow v_{\phi}(x,\hat{y}_{\delta})$triangleright Factual degradation of $\hat{y}_{\delta}$ in comparison to $x$
10:     Compute $\mathcal{M}_{\text{absolute}}\leftarrow\|\hat{\delta}_{\text{clean}}\|_{2}^{2}+\|\hat{\delta}_{\text{corrupt}}-\delta\|_{2}^{2}$$\triangleright$ Absolute consistency
11:     Compute $\mathcal{M}_{\text{relative}}\leftarrow\|\hat{\delta}_{\text{corrupt}}-\hat{\delta}_{\text{clean}}-\delta\|_{2}^{2}$$\triangleright$ Relative consistency
12:     $\mathcal{M}_{\text{MedVAL}}\leftarrow 1-\frac{1}{6}(\mathcal{M}_{\text{absolute}}+\mathcal{M}_{\text{relative}})$triangleright Generator-validator consistency score (0-1)
13:     if $\mathcal{M}_{\text{MedVAL}}\geq\tau$ then
14:         $\mathcal{D}_{\text{train}}\leftarrow\mathcal{D}_{\text{train}}\cup\{x,\hat{y},\hat{\delta}_{\text{clean}}\}$
15:         $\mathcal{D}_{\text{train}}\leftarrow\mathcal{D}_{\text{train}}\cup\{x,\hat{y}_{\delta},\hat{\delta}_{\text{corrupt}}\}$
16:     end if
17:end for
18:$v^{*}_{\alpha}=\textit{SFT}(v_{\alpha},\mathcal{D}_{\text{train}})$triangleright Supervised fine-tuning
19:return $v^{*}_{\alpha}$
Table 5: Language models included in our study for benchmarking and evaluation.
Language model Context Parameters Open-source? Medical adapted?
Llama 3.2 128k 3B
Qwen3 128k 4B
Llama 3.1 128k 8B
Gemma 3 128k 27B
MedGemma 128k 27B
Llama 3.3 128k 70B
GPT-4o Mini 128k unknown
GPT-4o 128k unknown
Claude Sonnet 4 200k unknown
Gemini 2.0 Flash 1000k unknown
Figure 5: Performance benchmark (Cohen’s $\kappa$). We rank all LMs (low to high). Baseline indicates zero-shot LM before fine-tuning, Baseline (larger comparator) indicates a larger zero-shot LM as reference (not chosen for fine-tuning), and MedVAL indicates LM after fine-tuning. Notably, smaller MedVAL LMs match or exceed the performance of much larger baselines.
Figure 6: MedVAL performance ablation. We report: a) the performance with self-distillation (for synthetic data generation) and without data filtering $\mathcal{M}_{\text{MedVAL}}\geq 0.0$, and b) the performance with self-distillation and data filtering $\mathcal{M}_{\text{MedVAL}}\geq 0.9$. The ablation confirms that MedVAL’s self-supervised data filtering positively contributes to LM performance improvements.
Figure 7: Representative examples of validation of LM-generated medical text by 1) the physician, 2) baseline GPT-4o, and 3) MedVAL fine-tuned GPT-4o. Under each example, MedVAL demonstrates higher agreement with the physician.
Table 6: Task-wise standard deviation (F1 score). Bootstrapped standard deviations reflect variability across predictions.
In-Distribution Out-of-Distribution
Open-Source
Baseline Llama-3.2-3B 0.011 0.014 0.022 0.016 0.033 0.039 0.011
Qwen3-4B 0.032 0.046 0.034 0.035 0.050 0.055 0.015
Llama3.1-8B 0.042 0.037 0.033 0.024 0.028 0.030 0.013
Gemma3-27B 0.031 0.036 0.036 0.027 0.041 0.054 0.014
MedGemma-27B 0.042 0.041 0.034 0.042 0.034 0.050 0.015
Llama-3.3-70B 0.032 0.047 0.035 0.038 0.039 0.056 0.017
MedVAL Llama-3.2-3B 0.040 0.033 0.036 0.034 0.037 0.047 0.017
Qwen3-4B 0.047 0.045 0.032 0.032 0.048 0.046 0.017
Llama-3.1-8B 0.046 0.048 0.029 0.030 0.042 0.045 0.017
Proprietary
Baseline GPT-4o Mini 0.034 0.045 0.035 0.032 0.033 0.056 0.015
GPT-4o 0.042 0.042 0.036 0.035 0.033 0.054 0.017
Claude Sonnet 4 0.041 0.051 0.035 0.035 0.048 0.057 0.017
Gemini 2.0 Flash 0.032 0.047 0.034 0.030 0.045 0.048 0.016
MedVAL GPT-4o Mini 0.039 0.028 0.032 0.036 0.041 0.063 0.018
GPT-4o 0.047 0.031 0.031 0.033 0.039 0.059 0.018

### v4

@@ 3.5 MedVAL Performance Ablation
Our ablation (Supplementary Figure 6) investigates the source of MedVAL’s improvements, demonstrating that gains arise from the "combined effect" of frontier LM distillation and MedVAL’s data filtering step. We evaluate two settings with no filtering ($\mathcal{M}_{\text{MedVAL}}\geq 0.0$) vs MedVAL filtering ($\mathcal{M}_{\text{MedVAL}}\geq 0.9$): 1) self-distillation (the model teaches itself), and 2) GPT-4o distillation (a frontier teacher).
We evaluate whether MedVAL helps even without a frontier teacher. Across models, when training on the same LM’s self-generated data, filtering consistently beats no filtering, and the self-distilled+MedVAL models also surpass their baselines. Notably, Llama-3.2-3B jumps from 12.8% $\to$ 22.1% F1 while using only 3% of the training set, whereas self-distilled without filtering moves just 12.8% $\to$ 13.6% despite using 100% of the data. This underscores that a small, high-quality subset can outperform far larger unfiltered corpora. We also see meaningful gains for Qwen3-4B (42.8% $\to$ 43.3%), Llama-3.1-8B (25.9% $\to$ 29.1%), and GPT-4o Mini (47.4% $\to$ 50.4%), indicating that a portion of the best GPT-4o-distilled performance is recovered by self-distilled+MedVAL, and that filtering drives improvements.
Using GPT-4o as a teacher improves F1 even without filtering, as expected; adding MedVAL on top of distillation yields the largest gains while being data-efficient. With MedVAL filtering, we fine-tune on 57% examples (1,131/2,000) yet still outperform the unfiltered 100% training set, for example, Llama-3.2-3B (33.4% $\to$ 42.9%), Qwen3-4B (51.1% $\to$ 52.7%), Llama-3.1-8B (43.8% $\to$ 46.5%), and GPT-4o Mini (49.5% $\to$ 54.0%). GPT-4o Mini, self-distilled+MedVAL (50.4%) even beats vanilla GPT-4o distilled without MedVAL (49.5%), showing MedVAL’s value beyond teacher strength. These results confirm that MedVAL is a data-efficient distillation method that can outperform vanilla distillation across student/teacher model variants.
3.6 External Validation under Distribution Shift
To probe robustness beyond our risk-grading benchmark, we evaluate MedVAL-distilled models on MEDEC [71], a benchmark for medical error detection/correction. Despite task and label differences, MedVAL improves accuracy over zero-shot baselines: from 53.3% $\to$ 54.4% for GPT-4o Mini and from 58.0% $\to$ 63.3% for GPT-4o. We choose GPT-4o and GPT-4o Mini models for this experiment as their baseline performance is reported on the MedHELM [72] leaderboard, enabling a fair, apples-to-apples comparison.
3.7 Comparison with Prior Metrics
We compare MedVAL with two representative metrics from Supplementary Table 4: 1) AlignScore [31], the only metric compatible with a physician-free, reference-free, retrieval-free setup, and 2) RadGraph F1 [44], a radiology-specific metric that requires reference outputs to compute overlap. We compute Pearson correlations between each metric and physician risk grades on the report2impression subset (n=190; a radiology task containing reference outputs). As shown in Supplementary Figure 10, MedVAL correlates strongly with physicians (GPT-4o: $r=0.825$, Qwen3-4B: $r=0.833$), AlignScore is moderate ($r=0.678$), and RadGraph F1 is weak ($r=0.156$). These findings support that reference-based similarity is not a faithful proxy for reference-free, input-only clinical risk grading, whereas MedVAL aligns closely with the physician rubric.
@@ 4 Discussion
We introduce MedVAL, a novel, self-supervised, and data-efficient distillation method for training LMs to validate LM-generated medical text following a physician-defined taxonomy. By combining synthetic data generation, data filtering, and fine-tuning, MedVAL improves a LM’s baseline capacity across diverse medical tasks. We validate MedVAL on MedVAL-Bench, a benchmark of 840 physician-annotated examples spanning 6 medical tasks. MedVAL extends prior methods by providing a framework that: 1) trains scalable evaluators without physician-in-loop supervision, 2) assesses medical text in the absence of reference outputs or retrieval, 3) supports multilingual evaluation, and 4) offers interpretable, expert-aligned assessments.
@@
Across all settings, MedVAL improved average four-class baseline F1 scores from 36.7% to 51.0%, with MedVAL GPT-4o achieving the highest score. MedVAL also improved average safe/unsafe binary F1 scores from 66.2% to 82.8%, indicating reliable discrimination between safe and unsafe responses. Risk-level analysis revealed that MedVAL enhances model sensitivity, particularly at intermediate risk levels (2–3), which are critical for deciding human review. Task-wise results confirmed strong generalization across in-distribution and out-of-distribution settings, with up to 84% relative improvement on held-out tasks. Notably, MedVAL displayed strong improvements on all unseen tasks (impression2simplified, bhc2spanish, dialogue2note) and an external benchmark (MEDEC), demonstrating robustness on challenging, real-world medical tasks.
@@
Similar to MedVAL, there exist automated methods for evaluating LM-generated medical text that have surfaced. However, these methods are either not medical domain-specific or require physicians-in-the-loop, the availability of reference outputs, or external knowledge bases for retrieval. MedVAL does not require either of these constraints. FActScore [30] uses domain-specific extractors for checking claims, though it relies on structured knowledge bases. AlignScore [31] proposes a unified alignment function for factuality assessment across tasks such as NLI, QA, and summarization. These methods, including model confidence calibration approaches [32, 33, 34], often lack the necessary nuance for clinically focused error assessment. In medical settings, there is often a continuum of right and wrong that requires a very nuanced understanding, making it uniquely challenging compared to general domains. Several studies have focused on error detection for LM-generated medical text. MedHAL [35], a benchmark for hallucination detection in medical text, relies on physician error annotations limited to specific curated medical tasks. DocLens [36] introduces multi-aspect metrics—completeness, conciseness, and attribution—tailored for medical text generation. While they report higher agreement with physician assessments than existing metrics, their evaluation relies on the availability of reference outputs. VeriFact [37] uses retrieval-based evidence to verify statements and is oriented towards multi-document summarization without the capability to train LMs. Recent studies have also focused on error assessment for text generation in radiology: ReXTrust [38] presents a fine-grained hallucination detector, ReXErr [39] injects clinical errors into radiology reports, GREEN [40] identifies significant errors in radiology reports, and FineRadScore [41] evaluates radiology report generation. While these efforts are crucial for identifying errors, their focus on radiology tasks (e.g., chest X-rays [42, 43, 44]) and reliance on reference outputs limits their generalizability. A detailed comparison of these methods is summarized in Supplementary Table 4.
Beyond methodological differences, real-world impact hinges on cost, latency, and privacy constraints. Hence, improving small/open models is directly relevant for real-world deployment as hospital systems face barriers to using proprietary APIs at scale. In agentic workflows, validation is a high-frequency step (per section, per note, per agent action), so efficient models are a practical path for routine validation, with frontier models reserved for escalations [73]. Furthermore, prior work has shown that modest-sized judges can track stronger teachers effectively [74] and on-policy self-judgment provides a privacy-preserving, label-efficient route to reliable evaluators [75]. MedVAL directly addresses these constraints by enabling data-efficient distillation, enabling small/open models to closely track frontier performance at a fraction of the cost and latency.
5 Limitations
Despite MedVAL’s broad applicability, our study has limitations. Our prompts for output-generation and perturbation injection were manually designed, potentially yielding partly “simulated” perturbations. Our goal was to enrich the test set with error families commonly observed by physicians [76], not replace natural errors. Furthermore, about one quarter of cases are unconstrained generations (Level 1), and even in Levels 2–4 models often deviate from instructions, yielding a blend of injected and naturally occurring errors. Nevertheless, we acknowledge that broader in-the-wild evaluations remain important future work.
@@ 5 Limitations
While our input-based validation approach enables reference-free evaluation, this can limit effectiveness in tasks where the input lacks sufficient information, such as question-answering (medication2answer). However, if reference outputs are available, they can be provided as an input to MedVAL, highlighting a beneficial extension. Furthermore, for QA tasks, we aimed to explore whether safety-critical decisions can be performed, regardless of whether the knowledge originates from the input or the LM’s knowledge base. Our framework demonstrated improved alignment with physicians in the medication2answer task, suggesting that MedVAL can enhance a LM’s ability to assess outputs leveraging its knowledge base.
Our task scope is not exhaustive. We evaluate six medical tasks that directly support risk-based triage. We acknowledge that this suite does not encompass the full spectrum of clinical document types and scenarios, such as EHR tabular reasoning, complex free-text reasoning, and consultation dialogues. Expanding MedVAL-Bench to these settings requires additional physician adjudication under our risk schema, which is a practical bottleneck; we therefore leave broader evaluation to future work, including representative tasks from MedHELM [72], MedS-Bench [77], and HealthBench [78]. Another limitation is that we use a single-pass consistency filter to select high-confidence training pairs, limiting early-stage noise amplification. While distillation using one-shot filtered data consistently outperforms distillation using all data unfiltered, iterative refinement schemes where the validator improves progressively with the filtered dataset could further improve data selection. However, we leave a systematic study of such strategies to future work.
Furthermore, a reasonable concern regarding our evaluation is that improvements could stem from simply learning the perturbation distribution. However, while supervision is synthetic (teacher signals filtered by generator-validator consistency), all evaluation uses physician labels. Moreover, gains persist on held-out tasks excluded from distillation (impression2simplified, bhc2spanish, and dialogue2note), which introduce distribution shifts. Together, these results suggest generalization beyond the perturbation style. Regardless, we leave a controlled study isolating perturbation design effects from error generalization to future works. Additionally, inter-physician agreement in our study was estimated on a stratified multi-annotated subset (15 examples $\times$ 6 tasks; 90/840) due to physician time constraints. While this design provides per-task coverage, it is a sparse sample; reliability estimates using Krippendorff’s $\alpha$ should be interpreted with this caveat. Finally, while we selected representative open-source and proprietary models for fine-tuning under practical resource settings, results may differ with larger models or other closed platforms. However, we open-source our code and trained models to allow the community to validate on broader tasks and models of interest.
@@ 6 Conclusion
In this research, we introduce MedVAL, a novel, self-supervised distillation method for training LMs to validate LM-generated medical text. Our benchmark highlights substantial improvements in model performance and provides systematic evidence that LMs can achieve performance statistically non-inferior to a single human expert ($p<0.001$). The MedVAL-Bench dataset, with its diverse tasks and physician annotations, establishes a robust benchmark for validating LM-generated medical text. To support further research and adoption, we release: 1) Codebase, 2) MedVAL-Bench, and 3) MedVAL-4B (the best-performing open-source model). Together, these contributions lay a foundation for developing and deploying safer, more trustworthy tools to automate the process of validating LM-generated medical text, reducing physician burden. The findings from this study suggest that incorporating MedVAL into clinical workflows could enhance the reliability of medical text generation, thereby improving patient safety and clinical efficiency. Future research should explore prospective studies to validate the practical benefits of MedVAL in real-world settings, potentially leading to an integrated approach that can support healthcare professionals and enhance patient care.
@@ 7 Data and Code Availability
Datasets used in this study include MedicationQA [47], MeQSum [48], Open-i [49], MIMIC-IV [50], MIMIC-IV-BHC [51] and ACI-Bench [52, 53, 54]. All these datasets (except MIMIC) are fully open-source. The MIMIC-IV datasets require PhysioNet access and are currently undergoing review to be made available on PhysioNet. Further distribution of datasets is subject to the data sharing agreements stipulated by the original creators.